OpenStudy (anonymous):

Find the side length of a square with the given area. 144 cm² A. 12 cm B. 36 cm C. 72 cm D. 288 cm

OpenStudy (anonymous):

Sqrt 144

OpenStudy (anonymous):

Find the square root of 144. :)

OpenStudy (anonymous):

sqrt(144)

OpenStudy (anonymous):

Same thing: \(A = s^2\) Plug in the area: \(144 = s^2\) Now find the square root of 144: \(s = \sqrt{144}\)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Yes